# Developmentally-plausible Working Memory Shapes a Critical Period for Language Acquisition

### v1

@@ Abstract
Large language models exhibit general linguistic abilities but significantly differ from humans in their efficiency of language acquisition. This study proposes a method for integrating the developmental characteristics of working memory during the critical period, a stage when human language acquisition is particularly efficient, into language models. The proposed method introduces a mechanism that initially constrains working memory during the early stages of training and gradually relaxes this constraint in an exponential manner as learning progresses. Targeted syntactic evaluation shows that the proposed method outperforms conventional models without memory constraints or with static memory constraints. These findings not only provide new directions for designing data-efficient language models but also offer indirect evidence supporting the underlying mechanisms of the critical period hypothesis in human language acquisition.
@@ 1 Introduction
An important theoretical framework for understanding the efficiency of human language acquisition is the Critical Period Hypothesis Lenneberg (1967). This hypothesis posits that there is a specific period during which language can be acquired efficiently, and that this ability diminishes thereafter. Various studies, including cases of limited language exposure during childhood and age-related effects on second-language acquisition, support the existence of a critical period Fromkin et al. (1974); Curtiss (1977); Johnson and Newport (1989). However, the reasons why children acquire language more efficiently than adults remain partially unresolved. One compelling explanation for the critical period in language acquisition is the Less-is-More Hypothesis Newport (1990), which argues that children’s cognitive limitations (e.g., working memory capacity and attentional scope) are advantageous for language learning. According to this hypothesis, children’s limited processing capacities enable them to efficiently extract fundamental patterns and structures (e.g., grammatical rules) from linguistic input, whereas adults, with their greater cognitive capacities, are more likely to be distracted by complex information, thereby hindering rule acquisition.
@@
Inspired by the “Less-is-More” hypothesis, this study investigates whether integrating human cognitive developmental characteristics, particularly the developmental properties of working memory (Figure 1), into language models can facilitate efficient language acquisition. Specifically, we propose a method for incorporating the exponential increase in working memory capacity that corresponds to the critical period into language models and analyze its impact on learning efficiency (§3). Using a GPT-2 model Radford et al. (2019) trained on a Child-Directed Speech (CDS) dataset Huebner and Willits (2021), we conduct evaluation experiments with Zorro Huebner et al. (2021), a targeted syntactic evaluation benchmark specialized for CDS (§4). The results demonstrate that a cognitively plausible model, which initially restricts working memory and gradually relaxes this constraint exponentially as training progresses, outperforms models without memory constraints or with static memory constraints. These findings provide new insights into designing data-efficient language models, contributing to the field of natural language processing, while also offering indirect evidence for the critical period hypothesis in human language acquisition, contributing to the field of cognitive science.
@@ 2 Related Work
2.1 Critical Period Hypothesis in Language Acquisition
The critical period hypothesis posits that language acquisition progresses most efficiently during a specific period, after which this ability declines. In the field of psycholinguistics, it has been noted that proficiency in pronunciation and grammar in second language acquisition (L2) declines sharply after adolescence, and this phenomenon is considered evidence supporting the existence of a biological critical period Tahta et al. (1981); Patkowski (1980). Furthermore, in the field of neuroscience, it has been suggested that the decline in brain plasticity with development and changes in neurotransmitter levels may contribute to the closure of the critical period Lenneberg (1967).
Nativist perspectives argue that language acquisition is a genetically programmed ability, with the critical period serving as its biological foundation Chomsky (1965); Pinker (1994). In contrast, empiricist perspectives propose that the effects of the critical period can also be explained by social and environmental factors, particularly the quantity and quality of linguistic input Elman et al. (1996); Seidenberg and Plaut (2006). Although numerous studies support the existence of a critical period, the scope of its impact on language ability, its precise endpoint, and the relative contributions of biological mechanisms and environmental factors remain subjects of ongoing debate.
@@ 2.2 The Role of Language Models in Acquisition Theories
In recent years, computational models have played a crucial role in elucidating the mechanisms of language acquisition. These models enable controlled investigations of learning mechanisms and environments, which are difficult to achieve with human participants, and they are used to test theoretical claims such as the “poverty of the stimulus” Clark and Lappin (2011). For instance, McCoy et al. (2020), Wilcox et al. (2024), and Warstadt et al. (2023) have employed language models to directly test hypotheses about language acquisition, demonstrating that such models can provide proof-of-concept evidence for learnability. These studies have attracted attention as efforts to deepen theoretical discussions on language acquisition through computational modeling, including research on the critical period for language acquisition.
Constantinescu et al. (2024) conducted experiments focusing on critical period phenomena in L2 acquisition [^1] The phenomenon in which delayed exposure to L2 makes learning more difficult and reduces proficiency. and first-language (L1) attrition. [^2] The phenomenon in which earlier cessation of L1 exposure increases the likelihood of forgetting L1. They found that, unlike human learners, language models do not naturally exhibit critical period effects. Furthermore, by incorporating a regularization method commonly used to mitigate catastrophic forgetting—Elastic Weight Consolidation Kirkpatrick et al. (2017)—they demonstrated that critical period effects could be artificially induced, mimicking the plasticity of the human brain. These results suggest that critical period effects are not an inevitable consequence of statistical learning but may instead be influenced by innate factors. While Constantinescu et al.’s (2024) research aimed to reproduce critical period effects in L2 acquisition and L1 attrition, this study focuses on reproducing critical period effects in L1 acquisition, highlighting a key distinction in objectives.
@@ 3.1 Modeling Developmental Trajectory of Human Working Memory
Based on these observations, the growth trajectory of working memory, as illustrated in Figure 1, is best characterized by an exponential model of the form $y=b-a^{x}$ ($10<a<$ 1). In this model, $b$ represents the asymptotic upper limit of working memory capacity, corresponding to adult-level performance, while $a$ determines the rate of growth. Specifically, smaller values of $a$ result in steeper early growth, reflecting the rapid cognitive development observed during early childhood, whereas larger values of $a$ indicate a slower rate of change.
@@ 3.2 Integrating Human Working Memory into Language Models
Here, $q_{i}\in\mathbb{R}^{1\times d}$, $K\in\mathbb{R}^{L\times d}$, $m$, and $B\in\mathbb{R}^{1\times L}$ represent the query, the key, a scalar slope specific to each attention head, and a bias matrix encoding the relative distances between queries and keys, respectively, where $B_{i,j}$ is defined as the negative absolute difference between positions $i$ and $j$.
The values of $m$ are set geometrically for each head. For example, in an 8-head model, the values of $m$ are assigned as follows: $m=1,\frac{1}{2},\frac{1}{4},\ldots,\frac{1}{128}$. The slope $m$ takes values in the range $[0,1]$, ensuring a consistent interpretation of its influence on attention scores.
By penalizing attention scores for query-key pairs with greater distances, ALiBi introduces a recency bias to the model. Originally, ALiBi was proposed to enhance the extrapolation capability of Transformer models. More recently, studies have shown that incorporating it into attention score computation during training allows for the estimation of surprisal patterns resembling human reading times Clark et al. (2025). This suggests its potential for modeling human-like memory decay and cognitive limitations.
However, since the slope $m$ in ALiBi is fixed for each attention head, the approach does not inherently reflect the developmental increase in working memory capacity (i.e., reduced decay) over time (Figure 1). Therefore, this study proposes a method, DynamicLimit-Exp, which replicates the developmental characteristics of working memory during the critical period, specifically its exponential growth. This is achieved by exponentially decreasing the slope $m$ in ALiBi as training epochs progress. In this method, the slope $m$ in the ALiBi mechanism is updated at each epoch $t$ as follows:
@@
Furthermore, a key distinction between ALiBi and DynamicLimit-Exp lies in how the slope $m$ is assigned across attention heads. While ALiBi applies a fixed per-head bias, enforcing a predetermined recency bias throughout training, DynamicLimit-Exp instead shares the slope $m$ across all heads. This ensures that the model maintains a globally coherent bias that evolves dynamically over the course of training. In other words, ALiBi imposes a head-specific static recency bias, whereas DynamicLimit-Exp introduces a dynamically changing proximity bias that governs the entire learning schedule. This shift enables the model to more accurately simulate the adaptive nature of human working memory development, potentially capturing critical periods of cognitive maturation.
@@ 4 Experiments
This study explores whether a language model trained from scratch can achieve more efficient language acquisition by incorporating the developmental characteristics of human working memory Specifically, we aim to determine whether this approach can replicate the increased efficiency of language acquisition observed during the critical period in L1 acquisition, focusing on the developmental advantages before the end of this period.
@@ 4.1 Configurations
Table 1: Accuracy (%) of models trained with AO-CHILDES. Overall represents the macro average of the scores across all grammar items.
@@
Table 2: Accuracy (%) of models trained with Wikipedia. Overall represents the macro average of the scores across all grammatical items.
For our experiments, we adopted GPT-2 Radford et al. (2019) as the base language model. While some studies utilize RoBERTa Liu et al. (2019) for base models Huebner et al. (2021); Warstadt et al. (2023), we selected GPT-2 for two primary reasons: (1) its unidirectional (left-to-right) predictions more effectively capture human short-term memory constraints (i.e., decay), and (2) GPT-based architectures dominate modern LLMs OpenAI (2023); Touvron et al. (2023).
@@
The training dataset, AO-CHILDES Huebner and Willits (2021), [^3] https://github.com/UIUCLearningLanguageLab/AOCHILDES is derived from the CHILDES dataset Macwhinney (2000), which records CDS from conversations between children and adults. AO-CHILDES contains 5 million words of speech directed at English-speaking children aged 1–6 years and controls for external factors such as age group, speaker variation, and situational context. As a preprocessing step, all sentences were converted to lowercase, and sentences shorter than three words were excluded. Since the AO-CHILDES dataset contains only about 5 million words, training a standard GPT-2 model would likely result in overfitting. To mitigate this, we followed existing studies on small language models (SLMs) trained with CDS datasets Huebner et al. (2021); Haga et al. (2024) and constructed an SLM with 4 layers, 4 attention heads, and 256 embedding dimensions for the base model. Details of the training configuration for the base model are provided in Appendix A.
Furthermore, to determine whether the critical period effect stems from exposure to specific linguistic stimuli, such as CDS, or from the model’s cognitive developmental properties independent of input, we conducted a complementary experiment using Wikipedia (written language, adult-oriented) as training data. Following Huebner et al. (2021), 500,000 sentences were randomly sampled from the English Wikipedia corpus. The latest version of Wikipedia, as of January 2025, was used as the dump file, [^4] https://dumps.wikimedia.org/enwiki/latest/enwiki-latest-pages-articles.xml.bz2 and preprocessing was performed using WikiExtractor. [^5] https://github.com/attardi/wikiextractor?tab=readme-ov-file
@@
We evaluate the grammatical abilities of these models using a developmentally inspired targeted syntactic evaluation benchmark, Zorro Huebner et al. (2021). Zorro is designed for assessing the syntactic and grammatical knowledge of language models in child-directed language and consists of 13 mid-level categories and 23 subcategories. Each subcategory contains 2,000 sentence pairs, with one grammatically acceptable and one unacceptable sentence per pair. Below is an example of a minimal pair from the “Subject-verb agreement” category: [^6] A full list of grammatical category in Zorro is provided in Appendix B
@@
By inputting both the acceptable and unacceptable sentence into the model and calculating the proportion of pairs where the model assigns a higher probability to the acceptable sentence, we obtain the grammaticality judgment score. In this study, we report scores for each mid-level category (henceforth, grammatical items) as well as their macro-average.

### v2

@@ Abstract
Large language models possess general linguistic abilities but acquire language less efficiently than humans. This study proposes a method for integrating the developmental characteristics of working memory during the critical period, a stage when human language acquisition is particularly efficient, into the training process of language models. The proposed method introduces a mechanism that initially constrains working memory during the early stages of training and gradually relaxes this constraint in an exponential manner as learning progresses. Targeted syntactic evaluation shows that the proposed method outperforms conventional methods without memory constraints or with static memory constraints. These findings not only provide new directions for designing data-efficient language models but also offer indirect evidence supporting the role of the developmental characteristics of working memory as the underlying mechanism of the critical period in language acquisition.
@@ 1 Introduction
An important theoretical framework for understanding the efficiency of human language acquisition is the Critical Period Hypothesis (CPH) Lenneberg (1967). The CPH posits that there is a specific period during which language can be acquired efficiently, and that this ability diminishes thereafter. Various studies, including cases of limited first language ($\mathrm{L}_{1}$) exposure during childhood and age-related effects on second language ($\mathrm{L}_{2}$) acquisition, support the existence of a critical period (CP) Fromkin et al. (1974); Curtiss (1977); Johnson and Newport (1989). However, the reasons why children acquire language more efficiently than adults remain partially unresolved. One compelling explanation for the CP in $\mathrm{L}_{1}$ acquisition is the Less-is-More Hypothesis Newport (1990), which argues that children’s cognitive limitations (e.g., working memory capacity and attentional scope) are advantageous for language learning. According to this hypothesis, children’s limited processing capacities enable them to efficiently extract fundamental patterns and structures (e.g., grammatical rules) from linguistic input, whereas adults, with their greater cognitive capacities, are more likely to be distracted by complex information, thereby hindering rule acquisition.
@@
Inspired by the “Less-is-More” hypothesis, we use language models (LMs) to study the CP for language acquisition, focusing on $\mathrm{L}_{1}$ acquisition and investigating whether integrating human cognitive developmental characteristics, particularly the developmental properties of working memory (Figure 1), into LMs can facilitate efficient language acquisition. Specifically, we propose a method for incorporating the exponential increase in working memory capacity that corresponds to the CP into LMs and analyze its impact on learning efficiency. Using a GPT-2 model Radford et al. (2019) trained on a Child-Directed Speech (CDS) dataset Huebner and Willits (2021), we conduct evaluation experiments with Zorro Huebner et al. (2021), a targeted syntactic evaluation benchmark specialized for CDS. The results demonstrate that a cognitively plausible model, which initially restricts working memory and gradually relaxes this constraint exponentially as training progresses, outperforms models without memory constraints or with static memory constraints. These findings provide new insights into designing data-efficient LMs, contributing to the field of natural language processing, while also offering indirect evidence supporting the role of the developmental characteristics of working memory as the underlying mechanism of the CPH in human language acquisition, contributing to the field of cognitive science.
@@ 2 Related Work
2.1 Critical Period for Language Acquisition
The CPH posits that language acquisition is most efficient within a specific developmental window, after which it declines. CP effects are observed in both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ acquisition, suggesting a shared underlying mechanism.
Critical Period for $\mathrm{L}_{1}$ Acquisition
Research in neurolinguistics and cognitive science suggests that there is a biologically determined CP for acquiring an $\mathrm{L}_{1}$, beyond which full native-like proficiency is unattainable if exposure to language is delayed. Studies on late $\mathrm{L}_{1}$ learners, such as deaf individuals who acquire sign language after early childhood, indicate severe deficits in grammatical proficiency compared to those exposed to language from birth Mayberry and Fischer (1989); Newport (1990). These findings suggest that neural plasticity, essential for $\mathrm{L}_{1}$ acquisition, diminishes with age, limiting the ability to develop full linguistic competence. From a theoretical perspective, the existence of the CP for $\mathrm{L}_{1}$ acquisition is often attributed to biological constraints. Nativist theories propose that $\mathrm{L}_{1}$ acquisition relies on an innate language faculty that operates most effectively during the CP Penfield (1965); Chomsky (1965); Pinker (1994). On the other hand, empiricist perspectives argue that the decline in $\mathrm{L}_{1}$ learning ability may result from environmental factors, such as a reduced need for language learning mechanisms once fundamental linguistic structures have been internalized Elman et al. (1996); Seidenberg and Zevin (2006). Despite extensive research, the precise boundary and mechanisms of the CP for $\mathrm{L}_{1}$ remain a subject of debate.
Critical Period for $\mathrm{L}_{2}$ Acquisition
CP effects are also observed in $\mathrm{L}_{2}$ acquisition, where late learners struggle with pronunciation, morphology, and syntax Johnson and Newport (1989); Hartshorne et al. (2018). While biological constraints play a role, entrenchment—where prior exposure to $\mathrm{L}_{1}$ limits flexibility in learning new linguistic structures—is also a factor Ellis and Lambon Ralph (2000); Seidenberg and Zevin (2006). Although the CP for $\mathrm{L}_{2}$ acquisition is an important topic, this study focuses on the CP for $\mathrm{L}_{1}$ acquisition, since our goal is to design data-efficient LMs by exploring the mechanisms of CP in $\mathrm{L}_{1}$ acquisition.
@@ 2.2 The Role of Language Models in Acquisition Theories
In recent years, computational models have played a crucial role in elucidating the mechanisms of language acquisition. These models enable controlled investigations of learning mechanisms and environments, which are difficult to achieve with human participants, and they are used to test theoretical claims such as the “poverty of the stimulus” Clark and Lappin (2011). For instance, McCoy et al. (2020), Wilcox et al. (2024), and Warstadt et al. (2023) have employed LMs to directly test hypotheses about language acquisition, demonstrating that such models can provide proof-of-concept evidence for learnability. These studies have attracted attention as efforts to deepen theoretical discussions on language acquisition through computational modeling, including research on the CP.
Constantinescu et al. (2025) investigated CP phenomena in $\mathrm{L}_{2}$ acquisition and $\mathrm{L}_{1}$ attrition, [^1] The phenomenon in which earlier cessation of $\mathrm{L}_{1}$ exposure increases the likelihood of $\mathrm{L}_{1}$ forgetting. assuming a shared underlying mechanism for CP effects across $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$. They simulated $\mathrm{L}_{2}$ exposure at varying ages to examine how LMs differ from human learners, finding that LMs do not naturally exhibit CP effects. To artificially induce such effects, they employed Elastic Weight Consolidation Kirkpatrick et al. (2017), a regularization method for mitigating catastrophic forgetting, thereby mimicking a maturational decline in plasticity. Their findings suggest that CP effects are not an inevitable outcome of statistical learning but may instead involve innate mechanisms.
While this study shares the broader objective of enhancing the cognitive plausibility of LMs as models of human language acquisition, it differs from Constantinescu et al. (2025) in both focus and methodology. Rather than modeling CP effects through dataset manipulation or post-CP plasticity constraints, this study explicitly addresses the developmental processes unfolding during the CP itself. Specifically, we integrate a mechanism to simulate the progressive growth of working memory capacity throughout the CP, a factor considered crucial for $\mathrm{L}_{1}$ acquisition but previously unmodeled in LM-based research. By incorporating developmental constraints, this study aims to provide a more fine-grained computational model of early $\mathrm{L}_{1}$ acquisition and its cognitive underpinnings, advancing the developmental plausibility of LMs.
@@ 3.1 Modeling Developmental Trajectory of Human Working Memory
Based on these observations, we characterized the growth trajectory of working memory, as illustrated in Figure 1, using an exponential model of the form $y=b-a^{x}$ ($10<a<$ 1). In this model, $b$ represents the asymptotic upper limit of working memory capacity, corresponding to adult-level performance, while $a$ determines the rate of growth. Specifically, smaller values of $a$ result in steeper early growth, reflecting the rapid cognitive development observed during early childhood, whereas larger values of $a$ indicate a slower rate of change.
@@ 3.2 Integrating Human Working Memory into Language Models
Here, $q_{i}\in\mathbb{R}^{1\times d}$, $K\in\mathbb{R}^{L\times d}$, $m\in\mathbb{R}_{[0,1]}$, and $B\in\mathbb{R}^{1\times L}$ represent the query, the key, a scalar slope specific to each attention head, and a bias matrix encoding the relative distances between queries and keys, respectively, where $B_{i}$ is defined as the negative absolute difference between the query position $i$ and each key position. The values of $m$ are set geometrically for each head. For example, in an 8-head model, the values of $m$ are assigned as follows: $m=1,\frac{1}{2},\frac{1}{4},\ldots,\frac{1}{128}$. The slope $m$ takes values in the range $[0,1]$, ensuring a consistent interpretation of its influence on attention scores. By penalizing attention scores for query-key pairs with greater distances, ALiBi introduces a recency bias to the model. Originally, ALiBi was proposed to enhance the extrapolation capability of Transformer models. More recently, Clark et al. (2025) has shown that incorporating it into attention score computation during training allows for the estimation of surprisal patterns resembling human reading times. This suggests its potential for modeling human-like memory decay and cognitive limitations.
However, since the slope $m$ in ALiBi is fixed for each attention head, the approach does not inherently reflect the developmental increase in working memory capacity (i.e., reduced decay) over time (Figure 1). Therefore, this study proposes a method, DynamicLimit-Exp, which replicates the developmental characteristics of working memory during the CP, specifically its exponential growth. This is achieved by exponentially decreasing the slope $m$ in ALiBi as training epochs progress. In this method, the slope $m$ in the ALiBi mechanism is updated at each epoch $t$ as follows:
@@
Furthermore, a key distinction between ALiBi and DynamicLimit-Exp lies in how the slope $m$ is assigned across attention heads. While ALiBi applies a fixed per-head bias, enforcing a predetermined recency bias throughout training, DynamicLimit-Exp instead shares the slope $m$ across all heads. This ensures that the model maintains a globally coherent bias that evolves dynamically over the course of training. In other words, ALiBi imposes a head-specific static recency bias, whereas DynamicLimit-Exp introduces a dynamically changing proximity bias that governs the entire learning schedule. This shift enables the model to more accurately simulate the adaptive nature of human working memory development, potentially capturing the CP of cognitive maturation.
@@ 4 Experiments
This study explores whether LMs trained from scratch can achieve more efficient $\mathrm{L}_{1}$ acquisition by incorporating the developmental characteristics of human working memory. Specifically, we aim to determine whether this approach can replicate the increased efficiency of $\mathrm{L}_{1}$ acquisition observed during the CP in $\mathrm{L}_{1}$ acquisition, focusing on the developmental advantages before the end of this period.
@@ 4.1 Configurations
Table 1: Accuracy (%) of models trained on AO-CHILDES dataset. Overall represents the macro average of the scores across all grammar items.
@@
Table 2: Accuracy (%) of models trained on Wikipedia dataset. Overall represents the macro average of the scores across all grammatical items.
We used the transformers Wolf et al. (2020) implementation of the GPT-2 Radford et al. (2019) as the base LM. While some studies utilize RoBERTa Liu et al. (2019) as a base model Huebner et al. (2021); Warstadt et al. (2023), we selected GPT-2 for two primary reasons: (1) its unidirectional (left-to-right) predictions more effectively capture human working memory constraints, and (2) GPT-based architectures dominate modern LLMs OpenAI (2023); Touvron et al. (2023b).
@@
We used AO-CHILDES Huebner and Willits (2021) [^2] https://github.com/UIUCLearningLanguageLab/AOCHILDES as the training dataset, which is derived from the CHILDES dataset Macwhinney (2000) and records CDS from conversations between children and adults. AO-CHILDES contains 5 million words of speech directed at English-speaking children aged 1–6 years and controls for external factors such as age group, speaker variation, and situational context. As a preprocessing step, following Haga et al. (2024), all sentences were converted to lowercase, and sentences shorter than three words were excluded. Since the AO-CHILDES dataset contains only about 5 million words, training a standard GPT-2 model would likely result in overfitting. To mitigate this, we followed existing studies on small language models (SLMs) trained with CDS datasets Huebner et al. (2021); Haga et al. (2024) and constructed an SLM with 4 layers, 4 attention heads, and 256 embedding dimensions for the base model. Details of the training configuration for the base model are provided in Appendix A.
Furthermore, to determine whether the CP effect stems from exposure to specific linguistic stimuli, such as CDS, or from the model’s cognitive developmental properties independent of input, we conducted a complementary experiment using Wikipedia (written language, adult-oriented) as training data. Following Huebner et al. (2021), 500,000 sentences were randomly sampled from the English Wikipedia corpus. We used the latest version of Wikipedia, as of January 2025,  [^3] https://dumps.wikimedia.org/enwiki/latest/enwiki-latest-pages-articles.xml.bz2 and preprocessed it using WikiExtractor. [^4] https://github.com/attardi/wikiextractor
@@
We evaluate the grammatical abilities of these models using a developmentally inspired targeted syntactic evaluation benchmark, Zorro Huebner et al. (2021). Zorro is designed for assessing the syntactic and grammatical knowledge of LMs in child-directed language and consists of 13 mid-level categories and 23 subcategories. Each subcategory contains 2,000 sentence pairs, with one grammatically acceptable and one unacceptable sentence per pair. Below is an example of a minimal pair from the “Subject-verb agreement (S-v agr)” category: [^5] See Appendix B for the full list of grammatical categories.
@@
By inputting both the acceptable and unacceptable sentence into the model and calculating the proportion of pairs where the model assigns a higher probability to the acceptable sentence, we obtain the grammaticality judgment score (Accuracy). In this study, we report scores for each mid-level category (henceforth, grammatical items) as well as their macro-average.
Model Overall D-n agr S-v agr Ana. agr Arg. str Binding Case Ellipsis Filler. gap Irregular Island Local. atr Quantifiers NPI AO-CHILDES DynamicLimit-Exp (↑) 62.2 50.8 50.0 49.6 67.7 58.7 95.2 43.1 93.6 52.2 53.6 51.3 57.6 85.0 DynamicLimit-Exp (↓) 56.5 49.9 49.7 50.1 44.7 61.9 70.6 73.3 72.0 51.8 61.9 47.0 48.1 54.1 $\Delta$ (↑, ↓) 5.7 0.9 0.3 -0.5 23.0 -3.2 24.6 -30.1 21.6 0.4 -8.3 4.4 9.5 30.8 Wikipedia DynamicLimit-Exp (↑) 59.1 49.8 50.4 46.0 71.5 59.3 97.7 37.4 86.5 51.1 58.0 60.5 42.2 53.9 DynamicLimit-Exp (↓) 52.9 50.4 50.1 47.4 68.7 62.3 74.4 60.2 44.2 53.2 51.7 42.7 40.6 42.2 $\Delta$ (↑, ↓) 6.1 -0.6 0.3 -1.4 2.9 -3.0 23.3 -22.8 42.3 -2.2 6.3 17.8 1.7 11.7
Table 3: Performance difference when changing the direction of the cognitive constraints in DynamicLimit-Exp